# OTTER: Improving Zero-Shot Classification via Optimal Transport

### v1

@@ 7 Conclusion
While zero-shot models have been successful, pretraining with massively collected dataset yields artifacts that may harm downstream tasks. In this paper, we identify the bias in class balance, and provide a simple but powerful solution using Optimal Transport. Theoretically, we describe how OT can fix label distribution mismatch and its sensitivity to perturbations. Empirically, we validated OT can successfully improve zero-shot classification accuracy, mitigating label distribution mismatch in zero-shot models. We believe our method can expedite the deployment of zero-shot classification, reducing the necessity of finetuning for downstream tasks.
@@ Acknowledgments
The appendix contains glossary, algorithm details, proofs, and detailed experimental results. The glossary contains a convenient reminder of our terminology (Appendix A). Appendix B describes the relevant algorithms used in our experiments, including Prior Matching [35] and BBSE [33]. Appendix C provides the proofs of theorems that appeared in Section 4. Finally, we give more details and analysis of the experiments and provide additional experiment results in Appendix D
@@ Appendix A Glossary
$[x]_{+}$ Take the positive parts of $x$, i.e. $[x]_{+}:=x\mathbbm{1}[x>0]$
$[x]_{-}$ Take the negative parts of $x$, i.e. $[x]_{-}:=x\mathbbm{1}[x<0]$
@@ Appendix B Algorithm details
[35] proposed prior matching as a reweighting method for prompt-based classifiers to mitigate word bias — the distribution shift between pretrained language models’ word prior and the class priors in the downstream task. We use it as a zero-shot model adaptation method given a class balance estimation.
@@
Though this is a non-convex optimization problem, it can be solved using the standard optimization techniques — we used [36]. While this is equivalently effective with OTTER when properly optimized, we found that the temperature parameter and learning rate crucially affect the final accuracy, making it less ideal for the zeroshot adaptation. We used the grid search with the small validation set (10 samples per class) in each task to select hyperparameters. The hyperparameter ranges are as follows.
@@
Label shift adaptation methods [33, 6, 22] aims to estimate the class balance in the target distribution using the labeled source distribution data and the unlabeled target distribution data. We use Black Box Shift Estimation (BBSE) to estimate the class balance in the downstream task. Algorithm 2 describes the procedure. Note that the derivation of this algorithm depends on the label shift assumptions, thus the class balance estimation with the reference data or the synthetic data can be heavily biased.
@@ Appendix C Theory details
Suppose $\hat{y}^{OT}_{i}\neq\hat{y}^{ZS}_{i}$ for some $i\in[n]$. It implies $\sum_{i=1}^{n}-\log P_{\theta}(Y=\hat{y}^{OT}_{i}|X=x_{i})<\sum_{i=1}^{n}-\log P_{\theta}(Y=\hat{y}^{ZS}_{i}|X=x_{i})$. However, this is a contradiction since, for any $i\in[n]$, $\hat{y}^{ZS}_{i}=\arg\max_{j\in[K]}P_{\theta}(Y=j|X=x_{i})$, thus $-\log P_{\theta}(Y=\hat{y}^{ZS}_{i}|X=x_{i})\leq-\log P_{\theta}(Y=j|X=x_{i})$ for all $j\in[K]$. ∎
@@
$=\arg\min_{j\in[K]}-\log P_{t}(Y=j|X=x_{i}),$
optimal transport solution
@@
gives Bayes classifier predictions by Theorem 4.1.
@@
$\pi^{*}=\arg\min_{\gamma\in\Pi(\mu,\nu)}\left\langle\gamma,C\right\rangle=\arg\min_{\gamma\in\Pi(\mu,\nu)}\left\langle\gamma,\tilde{C}\right\rangle.$
@@
The proof of Theorem 4.3 relies on the following result of [53].
Lemma C.3 ([53], Corollary 3.1.).
@@
where $1\leq p\leq\infty$ and $\kappa$ is the Hoffmann constant determined by $p,G,g$. [28].
@@
$\leq\kappa\left(\left|\left|(\Delta_{\nu})\right|\right|_{2}+\left|\left|[vec(\Delta_{C})]_{+}\right|\right|_{2}+\left|\left|vec(C)^{T}vec(\hat{\pi})-g^{T}\hat{w}\right|\right|_{2}\right).$
We adapt the definition for the Hoffman constant from [53]. Computing Hoffman constant or even bounding it has been a long-standing problem [5, 44, 45]. However, it has been shown that the Hoffman constant is a finite real number [53], and specifically under our problem setup, it is independent from perturbations and only related to original optimization problem. This suggests the possibility to regularize the parameters $C,G,g$ in the original problem such that $\kappa$ does not depend on the dimensionality of cost matrix or target distribution.
@@ C.4 Bounding label distribution estimation errors in few-shot learning
In few-shot learning, we assume that a few labeled samples per class are given. They can be used for estimating label distribution in the target distribution using label shift estimation methods [33, 6, 22]. They comes with the sample complexity analysis under the label shift assumptions, which can be used to obtain bound the label distribution estimation errors.
@@
The proof of Lemma C.4 relies on the following result of [32].
@@ D.1 Datasets
We use CIFAR10, CIFAR100 [31], Caltech101 [20], Caltech256 [23], Food101 [8], STL10 [14], SUN397 [68], Flower102 [41], EuroSAT [26], Oxford-IIIT Pet [43], Stanford Cars [30], DTD [13], CUB [63], ImageNet [16], ImageNet-r [27], and ImageNet-Sketch [65].
@@
We use Amazon [40], Gender [18], CivilComments [7], and HateXplain [37].
@@ D.2 Zero-shot classification setup
For zero-shot image classification, we emply CLIP [48] models. We used “a photo of a [CLASS]" prompt. Scores are computed by $s_{\theta}(x_{i},j)=P_{\theta}(Y=j|X=x_{i})=\cfrac{\exp\left(\cos(f(x_{i}),g(y_{j}))/\tau\right)}{\sum_{j^{\prime}=1}^{K}\exp\left(\cos(f(x_{i}),g(y_{j^{\prime}}))/\tau\right)}$ for image $x_{i}$ regarding the label $j$, given the image encoder $f$, the text encoder $g$. Cost matrix is constructed by $C=[c_{ij}]_{i\in[n],j\in[K]}$, where $c_{ij}=-\log s_{\theta}(x_{i},j)$. We run 1 with the true class balance of the test dataset.
@@
We employ BERT and text-embedding-ada-002 sentence embeddings for text classification [50]. This process parallels the methodology used in image zero-shot classification — we compute the prediction scores from the cosine similarity and then construct the cost matrix with the negative log probabilities.
D.3 Detailed experiment results of Section 5.2.1
@@
For the ablation study on zero-shot models, we provide experimental results with varying architectures, given the exact prior. Table 8, 9 show the image zero-shot classification results, and Table 10 shows the text zero-shot classification results. While the performance gain varies, we can observe that OTTER is effective for the most cases.
@@
Flower102
RN50 Zero-shot 66.5 38.70 74.7 74.1 76.3 93.1 57.8 57.3
Prior Matching 76.4(±$(\pm 0)$ 41.1(±$(\pm 2.5)$ 44.3(±$(\pm 9.7)$ 9.8(±$(\pm 1.6)$ 77.5(±$(\pm 1.4)$ 94.7(±$(\pm 0)$ 12.3(±$(\pm 3.9)$ 50.0(±$(\pm 2.8)$
OTTER(Ours) 76.8 44.6 83.6 80.2 81.1 95.5 63.8 64.3
RN101 Zero-shot 79.1 46.3 81.6 77.7 80.9 96.5 55.6 61.1
Prior Matching 80.3(±$(\pm 0.1)$ 46.7(±$(\pm 1.3)$ 57.6(±$(\pm 14.2)$ 9.6(±$(\pm 2.4)$ 80.8(±$(\pm 3.7)$ 96.7(±$(\pm 6)$ 17.6(±$(\pm 5)$ 49.0(±$(\pm 20.9)$
OTTER(Ours) 81.1 50.9 88.8 83.6 84.4 97.2 63.8 67.7
ViT-B/32 Zero-shot 88.9 58.5 81.3 79.5 80.2 97.1 61.1 59.3
Prior Matching 89.7(±$(\pm 0)$ 58.1(±$(\pm 1.2)$ 54.7(±$(\pm 14.1)$ 9.5(±$(\pm 1.6)$ 80.0(±$(\pm 4.4)$ 97.4(±$(\pm 0)$ 11.3(±$(\pm 3.1)$ 51.5(±$(\pm 3.2)$
OTTER(Ours) 89.7 64.4 88.1 84.9 85.0 97.8 68.0 68.1
ViT-B/16 Zero-shot 88.3 63.9 81.0 81.1 85.6 98.0 62.8 63.9
Prior Matching 91.3(±$(\pm 0)$ 64.1(±$(\pm 2.7)$ 59.3(±$(\pm 15.4)$ 9.5(±$(\pm 1.5)$ 86.8(±$(\pm 3.1)$ 98.4(±$(\pm 0)$ 6.7(±$(\pm 1.6)$ 54.0(±$(\pm 14.1)$
OTTER(Ours) 91.7 67.9 88.7 87.0 89.9 98.6 70.4 70.8
ViT-L/14 Zero-shot 95.0 72.3 80.3 85.0 89.8 99.2 64.9 72.3
Prior Matching 95.2(±$(\pm 0)$ 73.5(±$(\pm 2.5)$ 68.5(±$(\pm 19.7)$ 9.4(±$(\pm 1.2)$ 90.4(±$(\pm 3.8)$ 99.3(±$(\pm 0)$ 26.5(±$(\pm 1.1)$ 66.5(±$(\pm 17.9)$
OTTER(Ours) 96.0 77.7 92.0 90.9 93.6 99.4 71.5 81.3
Table 8: CLIP Zero-shot image classification accuracy (%)
EUROSAT
Pet
CUB
Stanford
Country211
@@
RN50 Zero-shot 18.2 80.5 45.6 13.3 37.3 47.1 51.5 34.9 5.3
Prior Matching 31.5(±$(\pm 6)$ 77.4(±$(\pm 0.3)$ 27.5(±$(\pm 1.9)$ 13.1(±$(\pm 0.6)$ 41.6(±$(\pm 0.2)$ 36.(±$(\pm 0.1)$ 44.6(±$(\pm 0.1)$ 7.8(±$(\pm 1.8)$ 5.(±$(\pm 0)$
OTTER(Ours) 26.8 83.0 49.2 14.1 42.2 51.7 54.1 37.1 5.7
RN101 Zero-shot 33.8 80.2 52.8 14.8 37.3 50.1 53.4 41.4 6.5
Prior Matching 34.0(±$(\pm 5.9)$ 77.2(±$(\pm 0.4)$ 34.8(±$(\pm 2.9)$ 15.1(±$(\pm 0.7)$ 39.4(±$(\pm 0.1)$ 33.6(±$(\pm 0.1)$ 46.7(±$(\pm 0.1)$ 9.5(±$(\pm 3.1)$ 6.5(±$(\pm 0.1)$
OTTER(Ours) 33.0 84.2 55.2 16.0 40.7 52.7 56.0 43.7 7.4
ViT-B/32 Zero-shot 29.7 81.7 49.0 15.5 40.7 51.4 55.6 60.7 34.3
Prior Matching 41.2(±$(\pm 0.7)$ 77.5(±$(\pm 0.6)$ 31.5(±$(\pm 1.1)$ 14.8(±$(\pm 0.1)$ 42.9(±$(\pm 0.1)$ 36.3(±$(\pm 0.9)$ 48.3(±$(\pm 0.1)$ 11.7(±$(\pm 0.8)$ $(\pm 0.3)$
OTTER(Ours) 44.9 86.2 52.10 15.9 45.1 54.7 57.7 64.2 39.4
ViT-B/16 Zero-shot 32.9 83.8 55.7 19.8 39.0 55.3 60.3 68.4 39.9
Prior Matching 41.6(±$(\pm 0.8)$ 82.(±$(\pm 0.3)$ 39.8(±$(\pm 2.6)$ 19.(±$(\pm 0.1)$ 42.1(±$(\pm 0.1)$ 40.4(±$(\pm 0)$ 53.6(±$(\pm 0.1)$ 16.6(±$(\pm 3.5)$ 36.5(±$(\pm 0.4)$
OTTER(Ours) 42.0 88.8 59.7 21.2 44.4 58.4 62.9 71.8 44.7
ViT-L/14 Zero-shot 25.75 87.93 64.12 28.2 50.8 61.8 67.6 80.6 51.9
Prior Matching 59.2(±$(\pm 0.4)$ 84.(±$(\pm 0.5)$ 53.(±$(\pm 15.6)$ 27.5(±$(\pm 0.3)$ 51.5(±$(\pm 0)$ 43.9(±$(\pm 0.3)$ 62.3(±$(\pm 0.1)$ 17.6(±$(\pm 5.1)$ 47.2(±$(\pm 0.4)$
OTTER(Ours) 57.62 91.03 70.02 29.5 51.0 66.3 70.2 83.3 55.2
@@
BERT Zero-shot 74.0 ($\pm$ 0.0) 84.1 ($\pm$ 0) 48.4 ($\pm$ 0.0) 30.9 ($\pm$ 0.0)
Prior matching 58.8 ($\pm$ 46.4) 41.4 ($\pm$ 39.6) 57.2 ($\pm$ 37.7) 31.3 ($\pm$3.3)
OTTER(Ours) 91.7 ($\pm$ 0.0) 91.9 ($\pm$ 0.0) 81.4 ($\pm$ 0.0) 34.3 ($\pm$ 0.0)
Ada Zero-shot 72.3 ($\pm$ 0.0) 50.9 ($\pm$ 0.0) 56.2 ($\pm$ 0.0) 27.9 ($\pm$ 0.0)
Prior matching 58.8 ($\pm$ 43.7) 50.0 ($\pm$ 41.3) 55.2 ($\pm$ 35.5) 32.4 ($\pm$ 4.0)
OTTER(Ours) 97.0 ($\pm$ 0.0) 73.7 ($\pm$ 0.0) 82.0 ($\pm$ 0.0) 32.0 ($\pm$ 0.0)
@@
Recent studies have demonstrated the efficacy of enhancing prompts as a means to improve zero-shot models [71, 38]. In order to further illustrate the potential enhancements offered by OTTER beyond enhanced prompts, we reproduced Menon and Vondrick [38]’s approach (Classification by Description, CD), which employs multiple prompts generated by language models and takes max scores of them for each class. We applied OTTER to CD. The results of this experiment are summarized in Table 12. As anticipated, OTTER exhibits enhancements in zero-shot classification, even when prompt sets are refined using language models.

### v2

@@ 6 Conclusion
While zero-shot models have been successful, pretraining using Internet-scale datasets yields artifacts that may harm downstream tasks. In this paper, we identify the bias in class balance, and provide a simple but powerful solution using optimal transport. Theoretically, we describe how OT can fix label distribution mismatch and its sensitivity to perturbations. Empirically, we validated our approach’s ability to improve zero-shot classification accuracy, mitigating label distribution mismatch in zero-shot models. We believe our method can expedite the deployment of zero-shot classification, reducing the necessity of finetuning for downstream tasks.
@@ Acknowledgments
The appendix contains glossary, algorithm details, proofs, and detailed experimental results. The glossary contains a convenient reminder of our terminology (Appendix A). Appendix B describes the relevant algorithms used in our experiments, including Prior Matching [liusie2023mitigating] and BBSE [lipton2018detectingbbse]. Appendix C provides the proofs of theorems that appeared in Section 4. Finally, we give more details and analysis of the experiments and provide additional experiment results in Appendix D
@@ Appendix A Glossary
$[x]_{+}$ Positive parts of $x$, i.e. $[x]_{+}:=x\mathbbm{1}[x>0]$
$[x]_{-}$ Negative parts of $x$, i.e. $[x]_{-}:=x\mathbbm{1}[x<0]$
@@ Appendix B Algorithm details
[liusie2023mitigating] proposed prior matching as a reweighting method for prompt-based classifiers to mitigate word bias — the distribution shift between pretrained language models’ word prior and the class priors in the downstream task. We use it as a zero-shot model adaptation method given a class balance estimation.
@@
It can be solved using the standard optimization techniques — we used [loshchilov2018adamw]. While this is equivalently effective with OTTER when properly optimized, we found that the temperature parameter and learning rate crucially affect the final accuracy, making it less ideal for the zeroshot adaptation. We used the grid search with the small validation set (10 samples per class) in each task to select hyperparameters. The hyperparameter ranges are as follows.
@@
Label shift adaptation methods [lipton2018detectingbbse, azizzadenesheli2019regularizedrlls, garg2020unifiedmlls] aims to estimate the class balance in the target distribution using the labeled source distribution data and the unlabeled target distribution data. We use Black Box Shift Estimation (BBSE) to estimate the class balance in the downstream task. Algorithm 2 describes the procedure. Note that the derivation of this algorithm depends on the label shift assumptions, thus the label distribution estimation with real-world data can be heavily biased.
@@ Appendix C Theory details
Suppose $\hat{y}^{OT}_{i}\neq\hat{y}^{ZS}_{i}$ for some $i\in[n]$. It implies $\sum_{i=1}^{n}-\log P_{\theta}(Y=\hat{y}^{OT}_{i}|X=x_{i})<\sum_{i=1}^{n}-\log P_{\theta}(Y=\hat{y}^{ZS}_{i}|X=x_{i})$. However, this is a contradiction since, for any $i\in[n]$, $\hat{y}^{ZS}_{i}=\arg\max_{j\in[K]}P_{\theta}(Y=j|X=x_{i})$, thus $-\log P_{\theta}(Y=\hat{y}^{ZS}_{i}|X=x_{i})\leq-\log P_{\theta}(Y=j|X=x_{i})$ for all $j\in[K]$, which results in $\hat{y}^{OT}_{i}=\hat{y}^{ZS}_{i}$. ∎
@@
$=\arg\min_{j\in[K]}-\log P_{t}(Y=j|X=x_{i}).$
Then, optimal transport solution
@@
leads to Bayes classifier predictions by Theorem 4.1.
@@
$\pi^{*}=\arg\min_{\gamma\in\Pi(\mu,\nu)}\left\langle\gamma,C^{*}\right\rangle=\arg\min_{\gamma\in\Pi(\mu,\nu)}\left\langle\gamma,C\right\rangle.$
@@
The proof of Theorem 4.3 relies on the following result of [robinson1973bounds].
Lemma C.3 ([robinson1973bounds], Corollary 3.1.).
@@
where $1\leq p\leq\infty$ and $\kappa$ is the Hoffmann constant determined by $p,G,g$. [hoffman1952approximate].
@@
Then, we have
$\begin{Vmatrix}\begin{pmatrix}\bar{z}\\ \bar{w}\end{pmatrix}-\begin{pmatrix}\hat{z}\\ \hat{w}\end{pmatrix}\end{Vmatrix}_{2}^{2}\leq\kappa^{2}\begin{Vmatrix}[\hat{\nu}-\nu]\\ [p-\hat{p}]_{+}\\ p\hat{z}-g\hat{w}\end{Vmatrix}_{2}^{2}.$
Then,
$\|\bar{z}-\hat{z}\|_{2}^{2}\leq\kappa^{2}\begin{Vmatrix}[\hat{\nu}-\nu]\\ [p-\hat{p}]_{+}\\ p\hat{z}-g\hat{w}\end{Vmatrix}_{2}^{2}-\|\bar{w}-\hat{w}\|_{2}^{2}$
Plugging in $\bar{z}=vec(\pi),\hat{z}=vec(\hat{\pi}),\Delta_{\nu}=\hat{\nu}-\nu,\Delta_{C}=\hat{C}-C$, and rearranging, we obtain
$\|\pi-\hat{\pi}\|_{F}^{2}\leq\kappa^{2}\left(\|\Delta_{\nu}\|_{2}^{2}+\|[\text{vec}(\Delta_{C})]_{+}\|_{2}^{2}+\|\text{vec}(C)^{T}\text{vec}(\hat{\pi})-g^{T}\hat{w}\|_{2}^{2}\right)-\|w-\hat{w}\|_{2}^{2}.$
We use the definition of the Hoffman constant $\kappa$ from [robinson1973bounds]. Computing Hoffman constant or even bounding it has been a long-standing problem [aze2002sensitivity, pena2021new, pena2024easily]. However, it has been shown that the Hoffman constant is a finite real number [robinson1973bounds], and specifically under our problem setup, it is independent from perturbations and only related to original optimization problem. This suggests the possibility to regularize the parameters $C,G,g$ in the original problem such that $\kappa$ does not depend on the dimensionality of cost matrix or target distribution. We leave further exploration of $\kappa$ in this context as future work.
@@ C.4 Bounding label distribution estimation errors in few-shot learning
In few-shot learning, we assume that a few labeled samples per class are given. They can be used for estimating label distribution in the target distribution using label shift estimation methods [lipton2018detectingbbse, azizzadenesheli2019regularizedrlls, garg2020unifiedmlls]. They comes with the sample complexity analysis under the label shift assumptions, which can be used to obtain bound the label distribution estimation errors.
@@
The proof of Lemma C.4 relies on the following result of [lipton2018detecting].
@@ C.5 Bounding the perturbation on cost matrix
C.6 Proof of Theorem 4.4
We provide the proof of Theorem 4.4.
Proof.
By assumption, we have $P_{\theta}(Y|X)=P_{s}(Y|X)$. By the result of Theorem 4.2, $y_{\textsc{OTTER}}$ samples are generated by $\arg\max_{j\in[K]}P_{t}(Y=j|X=x)$. Suppose $r^{*}=P_{t}(Y)/P_{s}(Y)$. Then,
$P_{\theta,r^{*}}(Y=j|X=x)$ $=\cfrac{r_{j}^{*}P_{\theta}(Y=j|X=x)}{\sum_{j^{\prime}=1}^{K}r^{*}_{j^{\prime}}P_{\theta}(Y=j^{\prime}|X=x)}$
$=\cfrac{r_{j}^{*}P_{s}(Y=j|X=x)}{\sum_{j^{\prime}=1}^{K}r_{j^{\prime}}^{*}P_{s}(Y=j^{\prime}|X=x)}$
$=\cfrac{\frac{P_{t}(Y=j)}{P_{s}(Y=j)}\frac{P_{s}(X=x|Y=j)P_{s}(Y=j)}{P_{s}(X=x)}}{\sum_{j^{\prime}=1}^{K}\frac{P_{t}(Y=j^{\prime})}{P_{s}(Y=j^{\prime})}\frac{P_{s}(X=x|Y=j^{\prime})P_{s}(Y=j^{\prime})}{P_{s}(X=x)}}$
$=\cfrac{P_{t}(Y=j)P_{s}(X=x|Y=j)}{\sum_{j^{\prime}=1}^{K}P_{t}(Y=j^{\prime})P_{s}(X=x|Y=j^{\prime})}$
$=\cfrac{P_{t}(Y=j)P_{t}(X=x|Y=j)}{\sum_{j^{\prime}=1}^{K}P_{t}(Y=j^{\prime})P_{t}(X=x|Y=j^{\prime})}$
$=\cfrac{\frac{P_{t}(Y=j)P_{t}(Y=j|X=x)P_{t}(X=x)}{P_{t}(Y=j)}}{\sum_{j^{\prime}=1}^{K}\frac{P_{t}(Y=j^{\prime})P_{t}(Y=j^{\prime}|X=x)P_{t}(X=x)}{P_{t}(Y=j^{\prime})}}$
$=\cfrac{P_{t}(Y=j|X=x)}{\sum_{j^{\prime}=1}^{K}P_{t}(Y=j^{\prime}|X=x)}$
Thus, we have
$\ROTTER\displaystyle y_{\text{\ROTTER}}$ $=\arg\max_{j}P_{\theta,r^{*}}(Y=j|X=x)$
$=\arg\max_{j}\cfrac{P_{t}(Y=j|X=x)}{\sum_{j^{\prime}=1}^{K}P_{t}(Y=j^{\prime}|X=x)}$
$=\arg\max_{j}P_{t}(Y=j|X=x)$
$=y_{\textsc{OTTER}},$
which implies $r^{*}$ is an optimal parameter. ∎
@@ D.1 Datasets
We use CIFAR10, CIFAR100 [dataset_cifar], Caltech101 [dataset_caltech101], Caltech256 [dataset_caltech256], Food101 [dataset_food101], STL10 [dataset_stl10], SUN397 [dataset_sun397], Flower102 [dataset_flower102], EuroSAT [dataset_eurosat], Oxford-IIIT Pet [dataset_pet], Stanford Cars [dataset_stanfordcar], DTD [dataset_dtd], CUB [dataset_cub], ImageNet [dataset_imagenet], ImageNet-r [dataset_imagenet-r], and ImageNet-Sketch [dataset_imagenet-sketch].
@@
We use Amazon [amazon_data], Gender [dataset_gender_bias1], CivilComments [dataset_civilcomments], and HateXplain [mathew2021hatexplain].
@@ D.2 Zero-shot classification setup
For zero-shot image classification, we emply CLIP [CLIP_base] models. We used “a photo of a [CLASS]" prompt. Scores are computed by $s_{\theta}(x_{i},j)=P_{\theta}(Y=j|X=x_{i})=\cfrac{\exp\left(\cos(f(x_{i}),g(y_{j}))/\tau\right)}{\sum_{j^{\prime}=1}^{K}\exp\left(\cos(f(x_{i}),g(y_{j^{\prime}}))/\tau\right)}$ for image $x_{i}$ regarding the label $j$, given the image encoder $f$, the text encoder $g$. Cost matrix is constructed by $C=[C_{ij}]_{i\in[n],j\in[K]}$, where $c_{ij}=-\log s_{\theta}(x_{i},j)$. We run Algorithm 1 with the true class balance of the test dataset.
@@
We employ BERT and text-embedding-ada-002 sentence embeddings for text classification [bert_embedder]. This process parallels the methodology used in image zero-shot classification — we compute the prediction scores from the cosine similarity and then construct the cost matrix with the negative log probabilities.
D.3 Detailed experiment results of Section 5.1
@@
For the ablation study on zero-shot models, we provide experimental results with varying architectures, given the exact prior. Table 8, 9 show the image zero-shot classification results, and Table 10 shows the text zero-shot classification results. We also provide another baseline results with CLIPPR [kahana2022improvingclippr], which uses the label distribution for adapter training. CLIPPR is similar with Prior Matching in the point that it requires adapter training, but it has more adapter layers and additional loss function to make the predictions stick to the original prediction scores. While the performance gain varies, we can observe that OTTER is effective for the most cases.
@@
Country211
DTD
EUROSAT
Flower102
RN50 Zero-shot 66.5 38.7 73.5 73.0 13.3 37.5 18.2 57.4
CLIPPR 69.2 20.1 59.4 67.1 8.2 13.0 19.3 30.7
Prior Matching 76.4(±$(\pm 0)$ 41.1(±$(\pm 2.5)$ 44.3(±$(\pm 9.7)$ 9.8(±$(\pm 1.6)$ 13.11($\pm$0.6) 41.6($\pm$0.2) 31.5 (±6.0$\pm$ 6.0)± 6.0 ) 50.0($\pm$2.8)
OTTER(Ours) 76.8 44.4 83.7 80.5 14.1 42.2 26.8 64.3
RN101 Zero-shot 79.1 46.3 81.0 76.6 14.8 37.3 33.8 61.3
CLIPPR 80.4 24.1 52.5 65.4 7.6 5.9 37.1 30.5
Prior Matching 80.3 (±$(\pm 0.1)$ 46.7 (±$(\pm 1.3)$ 57.6 (±$(\pm 14.2)$ 9.6 (±$(\pm 2.4)$ 15.1 (±$(\pm 0.7)$ 39.4 (±$(\pm 0.1)$ 34.0 (±$(\pm 5.9)$ 49. (±$(\pm 20.9)$
OTTER(Ours) 81.1 50.9 89.4 83.7 16.0 40.7 33.0 67.7
ViT-B/32 Zero-shot 88.9 58.5 80.3 78.2 15.5 40.7 29.7 59.3
CLIPPR 89.9 37.8 55.5 69.3 8.5 8.4 31.8 29.8
Prior Matching 89.7 (±$(\pm 0)$ 58.1 (±$(\pm 1.2)$ 54.7 (±$(\pm 14.1)$ 9.5 (±$(\pm 1.6)$ 14.8 (±$(\pm 0.1)$ 42.9 (±$(\pm 0.1)$ 41.2 (±$(\pm 0.7)$ 51.5 (±$(\pm 3.2)$
OTTER(Ours) 89.7 64.4 88.0 85.0 15.9 45.1 44.9 68.1
ViT-B/16 Zero-shot 88.3 63.8 79.8 79.8 19.8 39.0 32.9 64.0
CLIPPR 90.3 41.6 60.3 75.4 12.0 9.1 36.4 36.2
Prior Matching 91.3 (±$(\pm 0)$ 64.1 (±$(\pm 2.7)$ 59.3 (±$(\pm 15.4)$ 9.5 (±$(\pm 1.5)$ 19.0 (±$(\pm 0.1)$ 42.1 (±$(\pm 0.1)$ 41.6 (±$(\pm 0.8)$ 54.0 (±$(\pm 14.1)$
OTTER(Ours) 91.7 67.9 88.7 87.0 21.1 44.4 42.0 70.8
ViT-L/14 Zero-shot 95.0 72.3 78.2 83.4 28.2 50.8 25.8 72.3
CLIPPR 96.2 57.9 66.8 80.2 17.4 12.8 31.5 44.3
Prior Matching 95.2 (±$(\pm 0)$ 73.5 (±$(\pm 2.5)$ 68.5 (±$(\pm 19.7)$ 9.4 (±$(\pm 1.2)$ 27.5 (±$(\pm 0.3)$ 51.5 (±$(\pm 0)$ 59.2 (±$(\pm 0.4)$ 66.5 (±$(\pm 17.9)$
OTTER(Ours) 96.0 77.7 91.7 90.7 29.5 51.0 57.6 81.4
Table 8: CLIP Zero-shot image classification accuracy (%)
@@
RN50 Zero-shot 76.3 80.5 45.6 93.1 42.4 39.8 51.5 35.0 5.3
CLIPPR 69.5 61.9 32.5 94.4 45.0 19.0 31.1 23.8 0.7
Prior Matching 77.5 (±$(\pm 1.4)$ 77.4 (±$(\pm 0.3)$ 27.5 (±$(\pm 1.9)$ 94.7 (±$(\pm 0)$ 12.3 (±$(\pm 3.9)$ 36. (±$(\pm 0.1)$ 44.6 (±$(\pm 0.1)$ 7.8 (±$(\pm 1.8)$ 5.0 (±$(\pm 0.0)$
OTTER(Ours) 81.1 82.9 49.2 95.5 48.1 42.7 54.0 37.6 5.8
RN101 Zero-shot 80.9 80.2 52.8 96.5 36.9 34.9 53.4 41.9 6.5
CLIPPR 72.3 57.7 31.9 97.2 40.2 9.6 25.7 26.9 0.8
Prior Matching 80.8 (±$(\pm 3.7)$ 77.2 (±$(\pm 0.4)$ 34.8 (±$(\pm 2.9)$ 96.7 (±$(\pm 6)$ 17.6 (±$(\pm 5)$ 33.6 (±$(\pm 0.1)$ 46.7 (±$(\pm 0.1)$ 9.5 (±$(\pm 3.1)$ 6.5 (±$(\pm 0.1)$
OTTER(Ours) 84.4 84.2 55.2 97.2 44.6 41.0 56.0 44.5 7.3
ViT-B/32 Zero-shot 80.2 81.7 49.1 97.1 39.1 39.9 55.6 61.0 34.2
CLIPPR 73.9 58.0 30.6 97.5 41.6 11.9 32.6 51.7 19.9
Prior Matching 80.0 (±$(\pm 4.4)$ 77.5 (±$(\pm 0.6)$ 31.5 (±$(\pm 1.1)$ 97.4 (±$(\pm 0.0)$ 11.3 (±$(\pm 3.1)$ 36.3 (±$(\pm 0.9)$ 48.3 (±$(\pm 0.1)$ 11.7 (±$(\pm 0.8)$ 30.4 (±$(\pm 0.3)$
OTTER(Ours) 85.0 86.2 52.1 97.8 44.7 45.4 57.7 64.7 39.3
ViT-B/16 Zero-shot 85.6 83.8 55.7 98.0 47.1 46.0 60.2 68.9 39.8
CLIPPR 80.5 64.5 40.7 98.6 50.4 20.5 37.7 59.8 25.6
Prior Matching 86.8 (±$(\pm 3.1)$ 82.0 (±$(\pm 0.3)$ 39.8 (±$(\pm 2.6)$ 98.4 (±$(\pm 0)$ 6.7 (±$(\pm 1.6)$ 40.4 (±$(\pm 0)$ 53.6 (±$(\pm 0.1)$ 16.6 (±$(\pm 3.5)$ 36.5 (±$(\pm 0.4)$
OTTER(Ours) 89.9 88.8 59.7 98.6 54.1 50.4 62.9 72.4 44.5
ViT-L/14 Zero-shot 89.8 87.9 64.1 99.2 64.9 47.3 67.5 80.9 51.9
CLIPPR 87.9 82.7 54.0 99.6 68.2 27.3 45.4 74.5 35.5
Prior Matching 90.4 (±$(\pm 3.8)$ 84.0 (±$(\pm 0.5)$ 53.0 (±$(\pm 15.6)$ 99.3 (±$(\pm 0)$ 26.5 (±$(\pm 1.1)$ 43.9 (±$(\pm 0.3)$ 62.3 (±$(\pm 0.1)$ 17.6 (±$(\pm 5.1)$ 47.2 (±$(\pm 0.4)$
OTTER(Ours) 93.6 91.0 70.0 99.4 71.5 55.4 70.2 83.7 55.2
@@
BERT Zero-shot 74.0 84.1 48.4 30.9
CLIPPR 74.9 84.2 49.4 31.0
Prior Matching 58.8 ($\pm$ 46.4) 41.4 ($\pm$ 39.6) 57.2 ($\pm$ 37.7) 31.3 (±3.3$\pm$ 3.3)± 3.3 )
OTTER(Ours) 91.7 91.9 81.4 34.3
Ada Zero-shot 72.3 50.9 56.2 27.9
CLIPPR 73.7 53.0 57.6 27.6
Prior Matching 58.8 ($\pm$ 43.7) 50.0 (±41.3$\pm$ 41.3)± 41.3 ) 55.2 ($\pm$35.5) 32.4 ($\pm$ 4.0)
OTTER(Ours) 97.0 73.7 82.0 32.0
@@
Recent studies have demonstrated the efficacy of enhancing prompts as a means to improve zero-shot models [zhou2022learnprompt, menon2022visual]. In order to further illustrate the potential enhancements offered by OTTER beyond enhanced prompts, we reproduced menon2022visual’s approach (Classification by Description, CD), which employs multiple prompts generated by language models and takes max scores of them for each class. We applied OTTER to CD. The results of this experiment are summarized in Table 12. As anticipated, OTTER exhibits enhancements in zero-shot classification, even when prompt sets are refined using language models.